The front wheels of a wagon measure 3.5ft in diameter. The rear wheels measure 4.25ft in diameter. When the wagon is stopped, a chalk mark is made on both a front and a rear wheel. How far must the wagon travel before both chalk marks return to their initial position at the same time?
Answer by ankor@dixie-net.com(22740)

The front wheels of a wagon measure 3.5ft in diameter. The rear wheels measure 4.25ft in diameter. When the wagon is stopped, a chalk mark is made on both a front and a rear wheel. How far must the wagon travel before both chalk marks return to their initial position at the same time?
:
Find the front wheel circumference:
C = pi*3.5 = 10.99557 ft
:
Find the rear wheel circumference:
C = pi*4.25 = 13.35177 ft
:
10.99557 * 13.35177 = 146.8 ft